OpenStudy (anonymous):

Math Check please :) Solve \[\sqrt{9x+55}=x+5\] I got (x+30)(x-1) so -30 and 1

OpenStudy (anonymous):

square both side 9x+55= x^2+ 10x+25 x^2+x-30=0 (x+6)(x-5)=0

OpenStudy (saifoo.khan):

9x + 55 = (x+5)^2

OpenStudy (anonymous):

ahh imran your answer makes a whole lot more sense than mine.

OpenStudy (anonymous):

don't forget to check your answers! only one will work i think